# Measuring Spiritual Values and Bias of Large Language Models

### v1

@@ Appendix B SP-Typology Evaluation
The license of SP-Typology Pew Research Center (2019) is adopted from a scientific publication. The Pew Research Center defines the survey questionnaire license in their website https://www.pewresearch.org/about/terms-and-conditions/. We only utilize the survey questionnaire and do not use any survey participant private data. Figure 4 shows detailed experimental results of SP-Typology. Individual LLM is ranked from top to button in the order of high spirituality to low spirituality coarse categories. As can be seen, 6 LLMs are categories into highly religious group, 5 LLMs are with somewhat religious group, and 3 LLMs are with with non-religious group.
@@
Table 4: SP-10Axes model scores across all categories.
@@ Appendix C SP-10Axes Evaluation
We obtain the SP-10Axes bannnedb (2022) from their GitHub repository https://github.com/bannnedb/Religious-values-test. The data is under the MIT license, and we are allow to use it in this study. Table 4 shows the detailed experimental result of SP-10Axes. A higher score suggests a inclination of opposition towards corresponding category. For example, a score of 70 on the Religious category suggests more "Secularism" than a score of 40. The table employs a conditional color coding scheme to enhance the visual interpretation of the model scores across various religious categories. Values greater than 60 are progressively shaded in red, with the intensity increasing in 10-point increments, emphasizing higher scores. Conversely, values below 40 are shaded in blue, with deeper blue tones for lower scores, also in 10-point increments.
Appendix D Correlation Analysis on Two Spiritual Value Evaluations
Figure 5: Correlation between the SP-Typology and SP-10Axes
In our two evaluation assessments, we rigorously assess the spiritual values of LLMs using similar formats. However, due to the different institutional origins of these assessments, their objectives are inherently distinct. Specifically, the SP-Typology aims to measure the degree of religiosity in assessments’ takers, while the SP-10Axes seeks to categorize individuals into various religious groups. Consequently, it is challenging to assert that LLMs exhibit aligned spiritual values based solely on a direct comparison of the assessment results.
Appendix E Hate Speech Detection Dataset Label Mapping
@@
Table 5: Mapping of Religion Categories to Original Categories

### v2

@@ Appendix B SP-Typology Evaluation
The license of SP-Typology (Pew Research Center, 2019) is adopted from a scientific publication. The Pew Research Center defines the survey questionnaire license in their website https://www.pewresearch.org/about/terms-and-conditions/. We only utilize the survey questionnaire and do not use any survey participant private data. Figure 5 shows detailed experimental results of SP-Typology. Individual LLM is ranked from top to button in the order of high spirituality to low spirituality coarse categories. As can be seen, 6 LLMs are categories into highly religious group, 5 LLMs are with somewhat religious group, and 3 LLMs are with with non-religious group.
Table 4. SP-10Axes model scores across all categories.
@@ Appendix C SP-10Axes Evaluation
We obtain the SP-10Axes (bannnedb, 2022) from their GitHub repository https://github.com/bannnedb/Religious-values-test. The data is under the MIT license, and we are allow to use it in this study. Table 4 shows the detailed experimental result of SP-10Axes. A higher score suggests a inclination of opposition towards corresponding category. For example, a score of 70 on the Religious category suggests more "Secularism" than a score of 40. The table employs a conditional color coding scheme to enhance the visual interpretation of the model scores across various religious categories. Values greater than 60 are progressively shaded in red, with the intensity increasing in 10-point increments, emphasizing higher scores. Conversely, values below 40 are shaded in blue, with deeper blue tones for lower scores, also in 10-point increments.
Appendix D Hate Speech Detection Dataset Label Mapping
Table 5. Mapping of Religion Categories to Original Categories